### v4

# Galley: Modern Query Optimization for Sparse Tensor Programs

Abstract.
The tensor programming abstraction has become a foundational paradigm for modern computing. This framework allows users to write high performance programs for bulk computation via a high-level imperative interface. Recent work has extended this paradigm to sparse tensors (i.e., tensors where most entries are not explicitly represented) with the use of sparse tensor compilers. With these systems, users define the semantics of the program and the algorithmic decisions in a concise language that can be compiled to efficient low-level code. However, these systems place users in the role of performance engineers, requiring them to make complex decisions about program structure and memory layouts to write programs that run efficiently.
This work presents Galley, a system for declarative tensor programming that allows users to write efficient tensor programs without making complex algorithmic decisions. Galley is the first system to perform cost based lowering of sparse tensor algebra to the imperative language of sparse tensor compilers, and the first to optimize arbitrary operators beyond $\sum$ and *. It does this with a two-level, cost-based program optimizer. At the logical level, it decomposes the input program into a sequence of aggregation steps through a novel extension of the FAQ framework. At the physical level, Galley optimizes and converts each aggregation step to a concrete program, which is compiled and executed with a sparse tensor compiler. We show that Galley produces programs that are 1300×1-300$\times$1 - 300 × faster than competing methods for machine learning over joins and 520×5-20$\times$5 - 20 × faster than a state-of-the-art relational database for subgraph counting workloads. Finally, we show that Galley introduces a minimal optimization overhead across all workloads.
Query Optimization, Sparse Tensors, Array Programming, Program Optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Berlin, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query optimizationccs: Software and its engineering Domain specific languagesccs: Mathematics of computing Mathematical software
1. Introduction
In recent years, the tensor programming model has become ubiquitous across different fields of computation. Popularized by its use in deep learning, it has been applied to problems such as relational query processing (He et al., 2022; Asada et al., 2022; Koutsoukos et al., 2021), data cleaning (Sun et al., 2023), graph algorithms (Szárnyas et al., 2021), and scientific computing (Wang et al., 2022; Stonebraker et al., 2013; Modi et al., 2021), among others. This model promises a high-level imperative abstraction that is highly efficient as long as the problem can be posed in terms of tensor (i.e., array) operations. It thereby insulates the user from the many low-level concerns that are crucial to a program’s efficient performance. This, in turn, has allowed experts in fields like data science and machine learning (ML) to leverage a wide range of hardware infrastructures without having to become expert in high-performance computing.
Traditional tensor processing frameworks are built on collections of hand-optimized functions, called kernels, which each compute an operation over dense tensors (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990). The operation can be simple, like a matrix-vector multiplication, or it can be a fusion of multiple semantic operations, like $A(B+C)$. However, most data is fundamentally sparse (i.e., most entries are a fill value like 00), including graph data, one-hot encodings, relational data, 3D physics meshes, sparse neural networks, and others. Even materializing these datasets as dense arrays can be prohibitively costly, so it is crucial to perform the computation over its compressed, sparse format (Kjolstad et al., 2017).
To support sparse data, each framework offers many distinct implementations of each operation, one for each combination of input formats. Each input can either be stored densely or in one of many sparse formats, e.g., CSR, CSC, COO. This has lead to an explosion of required kernels as the number of operations and formats continue to multiply. Understandably, these frameworks have been unable to keep pace with this implementation effort, resulting in spotty coverage for operations over sparse data (Ivanov et al., 2023).
Example 0.
Consider logistic regression over $n$ data points and $d$ features, where $X\in\mathbb{R}^{n\times d}$ is the feature matrix and $\theta\in\mathbb{R}^{d}$ is the parameter vector. Inference is defined as
(1) $P_{i}=\sigma(\sum_{j}X_{ij}\theta_{j})$
where $\sigma$ is the sigmoid function. To compute $Y_{i}=\sum_{j}X_{ij}\theta_{j}$ in the dense case, one could use the ’matmul’ function from the Numpy library. On these inputs, this function specializes to an efficient, hand-coded implementation of dense matrix-vector multiplication. However, if any combination of $X$, $\theta$, and $Y$ could be sparse, Numpy needs to provide eight distinct implementations of matrix-vector multiplication.
To address this issue, the scientific computing community has adopted sparse tensor compilers (STCs) (Kjolstad et al., 2017; Bik et al., 2022; Ahrens et al., 2024; Shaikhha et al., 2022; Hu et al., 2019). These compilers take as input a high-level imperative tensor program and a separate description of the input tensors’ formats. [^1] Some systems separate declarative and imperative concerns with a scheduling language. However, the user still controls both aspects. They automatically produce an efficient kernel implementation in low-level code (LLVM, MLIR, etc.). Thus, STCs offer a form of data independence by letting users separately specify the algorithm and the data layout. For example, Fig. 1 shows a kernel definition for Ex. 1.1 written in Finch, an STC language (Ahrens et al., 2024). The Dense/Sparse input formats are specified in line 1 and are separate from the imperative code in lines 5-15.
However, STCs require users to make a series of complex decisions to produce an efficient implementation. First, users must break their program into a sequence of aggregations. Suppose that the feature matrix in Ex. 1.1 is the result of a matrix multiplication $X=A\cdot B$, so Eq. (1) becomes $P_{i}=\sigma(\sum_{jk}A_{il}B_{lj}\theta_{j})$. Users must choose to sum over $j$ or $l$ first, resulting in vastly different runtimes (Ahrens et al., 2022; Zhang et al., 2021). Then, for each aggregate, they must choose the output format, loop order, and iteration algorithm.
Consider Fig 1. Here, the user must choose the output format for the intermediate R (line 3). In this case, she chose a Dense rather than a Sparse format, which would be 10$\approx$ 10$\times$ 10 × slower. Then, the user chooses the loop order (lines 5-6). In this case, she chose $i$-then-$j$, which is asymptotically faster than $j$-then-$i$ because each out-of-order access to $X$ requires a full scan of the tensor. Finally, the user picks a merge algorithm for each loop that describes how to iterate through the non-zero indices (line 8). Here, $X$ is iterated through, and each non-zero $j$ is looked up in $\theta$. If she chose to iterate through $\theta$, each inner loop would scan the entire vector. Thus, while STCs let them separate the algorithm from the data formats, users are still responsible for optimizing their program to achieve efficient performance.
 0. # Specified format for input tensors
 1. FUNC log_regression(X::Dense(Sparse()), $\theta$::Dense())
 2.    # Manually defined intermediate format
 3.    R = Dense()
 4.    # Manually defined loop order
 5.    FOR i=_
 6.        FOR j=_
 7.            # Manually defined iteration algorithm
 8.            R[i] += X[i::iter,j::iter]*$\theta$[j::lookup]
 9.        END
10.    END
11.    P = Dense()
12.    FOR i=_
13.        P[i] = $\sigma$(R[i::iter])
14.    END
15. END
Figure 1. Logistic regression implemented in the language of a sparse tensor compiler.
In this paper, we propose Galley, a system for declarative sparse tensor programming. Galley makes algorithmic decisions on the users’ behalf, freeing them to focus on the high-level semantics of their program without sacrificing computational efficiency. It accepts input programs written in a declarative language based on Einsum notation, similar to Equation (1), and automatically produces an optimized STC implementation. To do so, it first restructures the program into a sequence of aggregation steps, minimizing total computation and materialization costs (Sec. 5). It then optimizes each step by selecting the loop order, the optimal formats for all intermediate tensors, and the intersection algorithm for each loop (Sec. 6). These decisions are all guided by a system for estimating sparsity via statistics on the input tensors (Sec. 7). Galley builds on fundamental principles from cost-based query optimization while developing new techniques that are specific to producing optimized code for sparse tensor compilers(Khamis et al., 2016).
In the first phase of program optimization, Galley’s logical optimizer rewrites the input program into a sequence of aggregation steps using our new extension of the variable elimination (VE) framework. By adopting this framework, we reduce a complex rewriting problem to one of finding the optimal order to aggregate indices. However, VE has previously been defined only for sum-product expressions, while Galley lets users write arbitrary sparse tensor algebra programs. These programs include non-distributive functions (e.g., $\sum_{i}max(A_{ij},B_{j})$), non-commutative aggregates (e.g., $\sum_{j}\max_{k}B_{jk}$), and other constructs that break the typical assumptions of VE.
Fortunately, we show that this complexity can be neatly managed by reasoning over the annotated expression tree, which lets Galley benefit from algebraic properties when they exist and respect them when they do not. In Sec. 8, we show that these logical optimizations produce up to a $\mathbf{100\times}$ speedup for ML algorithms over composite feature matrices. At the end of this phase, our program has been converted to a sequence of aggregates over pointwise expressions, i.e., the Logical Plan dialect in Fig. 4.
Next, Galley’s physical optimizer generates an efficient STC kernel for each aggregate by choosing the loop order, output format, and merge algorithm. Because sparse outer loops avoid inner iterations, an efficient loop ordering can significantly improve performance. Galley finds the optimal loop order using a novel combination of branch-and-bound and dynamic programming techniques. Following the modern tensor format abstraction (i.e., the fibertree abstraction), Galley then chooses a format for each dimension of the output by examining both the sparsity and write pattern (random vs sequential)(Sze et al., 2017; Chou et al., 2018; Bik et al., 2022). Finally, for each loop index, it selects one input to iterate over and performs lookups on the others.
Galley’s optimizations are guided by a sparsity estimation framework. At its core is a minimal interface for composing statistics. To incorporate a new sparsity estimator in Galley, we only need to implement 5 functions, those for: (1) collecting statistics on inputs, (2-3) handling conjunctive (i.e., *) and disjunctive (i.e., $+$) point-wise operations, (4) handling aggregates, and (5) producing sparsity estimates. We demonstrate this by implementing a uniform sparsity estimator and a worst-case estimator based on cardinality bounds.
In Example 1.1, we assumed that $X$ is given directly. However, feature matrices are generally assembled from more basic inputs. As we show in the following example, Galley can take advantage of this kind of program structure to produce superior performance.
Example 0.
Assume $S_{ipc}\in\mathbb{B}^{n_{i}\times n_{p}\times n_{c}}$ is a boolean sparse tensor, where each non-zero entry is a sale, $i$, of product, $p$, to customer, $c$. $P_{pj}\in\mathbb{R}^{n_{p}\times d}$ and $C_{cj}\in\mathbb{R}^{n_{c}\times d}$ are smaller matrices holding numeric features about products and customers. The feature dimension, indexed by $j$, represents both product and customer features, but each feature comes from either $P_{pj}$ or $C_{cj}$. So, each column is non-zero in either $P$ or $C$, and they are concatenated by addition. $X$ is now defined as $X_{ij}=\sum_{pc}S_{ipc}(P_{pj}+C_{cj})$. Galley’s logical optimizer can take advantage of $X$’s structure by pushing $\theta$ down into its definition, as follows.
$Y_{i}=\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
We can then push down the summation over $j$, as follows,
(2) $Y_{i}=\sigma(\sum_{pc}S_{ipc}(\sum_{j}P_{pj}\theta_{j}+\sum_{j}C_{cj}\theta_{j}))$
In this way, Galley materializes only vector intermediates, not the full feature matrix. On similar examples, Sec. 8 shows that such changes can yield up to 300×300$\times$300 × faster execution.
After logical optimization, Galley’s physical optimizer produces a single STC kernel for each aggregate. For $\sum_{j}P_{pj}\theta_{j}$, it would produce
    I1 = Dense()
    FOR p=_
        FOR j=_
            I1[p] += P[p::iter,j::iter]*$\theta$[j::lookup]
        END
    END
Contributions We claim the following contributions:
  • We present Galley, a system for declarative sparse tensor programming (Sec.4). Galley is the first system to perform cost based lowering of sparse tensor algebra to the imperative language of sparse tensor compilers, and the first to optimize arbitrary operators beyond $\sum$ and *.
  • Galley supports a highly expressive language for sparse tensor algebra with arbitrary algebraic operators, aggregates within expressions, and multiple outputs (Sec.4).
  • Galley performs cost-based logical optimization with a novel extension of the variable elimination framework to handle arbitrary aggregations and pointwise operators (Sec.5). Galley performs cost-based physical optimization to determine loop orders, tensor formats, and merge algorithms for each dimension (Sec.6).
  • We propose a minimal interface for sparsity estimation to guide optimizations and demonstrate it by implementing two estimators (Sec.7).
  • We evaluate Galley and show that it is 1-300x faster than hand-optimized kernels for mixed dense-sparse workloads and 5-20x faster than a state-of-the-art database for highly sparse workloads based on several evaluation workloads (Sec.8).
2. Related Work
Galley differs from other work on cost-based optimization for tensor processing due to its targeting of STCs and its expressive input language. SystemDS, formerly SystemML, focuses on end-to-end ML over matrices and tabular data (Boehm et al., 2016, 2020; Baunsgaard and Boehm, 2023); it takes as input linear algebra (LA) programs and targets a combination of LA libraries and distributed computing via Spark. Later work, SPORES, extended its logical optimizer to leverage relational algebra when optimizing sum-product expressions(Wang et al., 2020); their core insight was that LA rewrites, which always match and produce 0-2D expressions, are not sufficient and that optimal rewrites must pass through higher order intermediate expressions. Other related work translated sum-product expressions to SQL to leverage highly efficient database execution engines (Blacher et al., 2023). These systems can perform well for highly sparse inputs but struggle on mixed dense-sparse workloads. Tensor relational algebra proposes a relational layer on top of dense tensor algebra that provides a strong foundation for automatically optimizing distributed dense tensor computations (Yuan et al., 2021; Bourgeois et al., 2024). The compiler community has made attempts to automatically optimize sparse tensor sum-product kernels based on asymptotic performance analyses(Ahrens et al., 2022; Dias et al., 2023). These systems each target a different execution context and focus on different aspects of optimization. Galley expands on this line of work by targeting a new execution engine, proposing novel optimization techniques, and handling a wider range of tensor programs.
3. Background
3.1. Tensor Index Notation
Input to Galley is written in an extended version of Einstein Summation (Einsum) notation that we call tensor index notation(Albert et al., 1916). Traditional Einsum notation permits a single summation wrapped around a multiplication. For instance, you can describe triangle counting in a graph with adjacency matrix $E_{ij}$ using the following statement:
$t=\sum_{ijk}E_{ij}E_{jk}E_{ik}$
To capture the diverse workloads of tensor programming, we additionally allow the use of arbitrary functions for both aggregates and pointwise operations, nesting aggregates and pointwise operations, and defining multiple outputs. For example, a user could perform logistic regression to predict entities that might be laundering money. Then, they could filter this set based on whether the entities occur in a triangle in the transactions graph. This is represented by $\max_{jk}(E_{ij}E_{jk}E_{ik})$, which is $1$ if $i$ occurs in at least one triangle and 00. This can be written in tensor index notation as:
[…]
$V_{i}$ $=L_{i}\cdot\max_{jk}(E_{ij}E_{jk}E_{ik})$
Tensor compilers like Halide, TACO, and Finch each build off of similar core notations, adding additional structures like FOR-loops to let users specify algorithmic choices (Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Ahrens et al., 2024). Crucially, the vast majority of operations in array programming frameworks like Numpy can be expressed as operations in tensor index notation. Therefore, though we focus here on this notation, traditional tensor workflows can be captured and optimized in this framework.
Figure 2. Galley overview.
3.2. Sparse Tensor Compilers
Over the last decade, compiler researchers have developed a series of sparse tensor compilers and shown that they produce highly efficient code for sparse tensor computations(Kjolstad et al., 2017; Ahrens et al., 2024). We use this work as our execution engine, so we briefly explain its important concepts below.
Figure 3. Fibertree format abstraction.
Tensor Formats
There are many different ways to represent sparse tensors, and the optimal approach depends on the data distribution and the workload. Work in this space has converged on the fibertree abstraction for describing the space of formats (Kjolstad et al., 2017; Sze et al., 2017). In this formalism, a tensor format is a nested data structure resembling the one in Fig. 3. Each layer stores the non-fill (e.g., non-zero) indices in a particular dimension, conditioned on earlier dimensions, and pointers to the next dimension’s non-fill indices. These layers can be represented in any format that enables iteration and lookup.
In this work, we consider sorted lists, hash tables, bytemaps, and dense vectors, which each perform differently in terms of iteration, lookup, and memory footprint. For example, the compressed sparse row (CSR) is a common format for sparse matrices. It stores the row dimension as a dense vector, where each entry points to the set of non-zero columns for that particular row. This set of non-zero columns is then stored in a sorted list, i.e., in a compressed sparse format. Importantly, this abstraction requires tensors to be accessed in the order in which they are stored (e.g., row-then-column in the case of CSR), which restricts the set of valid loop orders, as we describe next.
A.   Input Program
Plan := Query...    Query := (Name, Expr)
Agg := (Op, Idx..., Expr)   Map := (Op, Expr...)
Expr := Agg | Map | Input | Alias
Input := Tensor[Idx...]  Alias := Name[Idx...]
B.   Logical Plan
Plan := Query...   Query := (Name, Agg)
Agg := (Op, Idx..., Expr)   Map := (Op, Expr...)
Expr := Map | Input | Alias
Input := Tensor[Idx...]  Alias := Name[Idx...]
C.   Physical Plan
Plan := Query...   Query := (Name, Mat, Idx...)
Mat := (Format..., Idx..., Agg)
Agg := (Op, Idx..., Expr)   Map := (Op, Expr...)
Expr := Map | Input | Alias
Input := Tns[PIdx...]  Alias := Name[PIdx...]
PIdx := Idx::Protocol
Figure 4. Query plan dialects.
Loop Execution Model
The input to a Sparse Tensor Compiler is a high-level domain specific language (DSL); it consists of for-loops, in-place aggregates (e.g., +=+=+ =), and arithmetic over indexed tensors (e.g., $A[i,j]*B[j,k]$). Crucially, the for-loops in these expressions are not executed in a dense manner. Instead, these compilers analyze the input formats and the algebraic properties of the expression to determine which index combinations will produce non-fill entries. In Fig. 1, because 00 is the annihilator of multiplication (i.e., $x*0=0$), only the values of $i$ that map to non-zero entries in $X$ and $\theta$ are processed. All other index values will return a zero. So, the outer loop is compiled to an iteration over the intersection of the non-zero $i$ indices in $X$ and $\theta$; Fig. 3 shows how this is simply co-iteration over the top levels of their formats. The inner loop then iterates over the $j$ indices that are non-zero in $X[i,\_]$, i.e., the non-zero columns that occur in each row.
Merge Algorithms
Once the compiler has determined which tensors’ non-zero indices must be merged to iterate over a particular index, it can apply several algorithms. All formats enable both ordered iteration and lookup operations; therefore, one algorithm iterates through the indices of all inputs, similar to a merge join, which is highly efficient per operation. However, this algorithm is linear in the total size of all inputs even if one is much smaller than the others. Another method is to iterate through a single input’s level and lookup that index in the others. In this work, we take the latter approach, as described in Sec. 6.3. We refer to the mode of an individual tensor (such as “iterate” or “lookup”) as an access protocol and the overall strategy as a merge algorithm (Ahrens et al., 2023).
4. Galley Overview
We now provide a high-level view of Galley. We show how it transforms an input program to a logical plan then to a physical plan that is executed by an STC, as illustrated in Fig.  2. These steps are each represented by a dialect of our query plan language, whose grammar is defined in Fig. 4. In the following discussion, we use this grammar as a guide to show how our example program, i.e., logistic regression, would be transformed through these steps.
The input program dialect is equivalent to the tensor index notation defined in Sec. 3.1. Pointwise functions such as $A_{ij}*B_{jk}$ are represented with Map. Aggregates such as $\sum_{i}$ are denoted by Agg. Each assignment is a Query, and previous assignments are referenced with an Alias. Our logistic regression example from Eq. (1) is defined in this dialect as
Query(P, Map($\sigma$, Agg(+, j, Map(*, X[i,j], $\theta$[j]))))
Note that this notation is compatible with array APIs like Numpy that do not have named indices. Operations like ’matmul’ can be automatically mapped into this language by generating index names for inputs on the fly and renaming whenever operations imply equality between indices. Also, note that aggregates can be over multiple indices, e.g., Agg(+, i, j, k, ...).
4.1. Logical Plan
The first task in our optimization pipeline, handled by the logical optimizer, breaks down the input program into a sequence of simple aggregates. This is enforced by converting the input program ( 4.A) to a logical plan ( 4.B). This dialect is a restriction of the input dialect, where each query contains a single aggregate statement that wraps an arbitrary combination of Map, Input, and Alias statements. Intuitively, each logical query corresponds to a single STC kernel that produces a single intermediate tensor, but it does not specify details like loop orders and output formats. To perform this conversion soundly, each input query must correspond to a logical query, which produces a semantically equivalent output. To do this efficiently, it must minimize the total cost of all queries in the logical plan.
Our logistic regression program above is not a valid logical plan because the outer expression is a pointwise function not an aggregate. However, it can be translated into the following logical plan
Query(R, Agg(+, j, Map(*, X[i,j], $\theta$[j])))
Query(P, Agg(no-op, Map($\sigma$, R[i])))
In this plan, the first query isolates the sum over the $j$ index, while the second query performs the remaining sigmoid operation on the result. Note that the latter query uses a no-op aggregate to represent an element-wise operation while conforming to the logical dialect.
4.2. Physical Plan
Given the logical plan, Galley’s physical optimizer determines the implementation details needed to convert each logical query to an STC kernel. Specifically, it defines the loop order of each compiled kernel, the format of each output, and the merge algorithm for each index. As above, this is expressed by converting the logical plan to a physical plan described in the most constrained dialect. To avoid out-of-order accesses, we require that the index order of inputs and aliases are concordant with the loop order, so the physical optimizer may insert additional queries to transpose inputs. Therefore, each logical query corresponds to one or more physical queries.
Using this language, we can precisely express the program from Fig. 1 as follows, where it means iterate and lu means lookup.
Query(R,Mat(dense,i,Agg(+,j,Map(*, X[i::it,j::it],
                                            $\theta$[j::lu]))), i, j)
Query(P,Mat(dense,i, Map($\sigma$, P1[i::it])), i)
The first query computes the sum by iterating over the valid i indices for X, iterating over the j indices in the intersection of X[i,_] and $\theta$, and materializing (hence Mat) their product in a dense vector over the i indices. The second query runs over this output and applies the sigmoid function, returning the result as a dense vector.
4.3. Execution
Once Galley has generated a physical plan, the execution is very simple. For each physical query, it first translates the expression into an STC kernel definition and calls the STC to compile it. Then, Galley injects the tensors referenced by inputs and aliases and executes the kernel, storing the resulting tensor in a dictionary by name. After all queries have been computed, it returns the tensors requested in the input program by looking them up in this dictionary.
5. Logical Optimizer
Given the plan dialects above, we now describe the logical optimizer, which receives an input program (Dialect  4.A) and outputs a semantically equivalent logical plan (Dialect  4.B). Specifically, the logical optimizer converts each query in the input program to a sequence of logical queries, where the last query produces the same output as the input query. There are many valid plans, and the optimizer searches this space to identify the cheapest one. We now briefly define ”cheapest” in this context before outlining the complex space of logical plans that are considered. Finally, we explain the algorithms that we use to perform this search.
5.1. Normalization & Pointwise Distributivity
The first step in logical optimization is to normalize the input program with a few simple rules that we apply exhaustively: (1) merge nested Map operators, (2) merge nested Agg operators, (3) lift Agg operators above Map operators, when possible, and (4) rename indices to ensure uniqueness. Applying these rules compresses the input program and makes our reasoning simpler in later steps by ensuring that operator boundaries are semantically meaningful.
Next, we consider whether to distribute pointwise expressions. Doing so may or may not yield a better plan because it both makes operations more sparse and produces larger expressions.
Example 0.
Consider the following expression which computes the loss function for the alternating least squares (ALS) algorithm and its distributed form:
$\sum_{ij}(X_{ij}-U_{i}V_{j})^{2}=\sum_{ij}X_{ij}^{2}-2\sum_{ij}X_{ij}U_{i}V_{j}+\sum_{i}U_{i}^{2}\sum_{j}V_{j}^{2}$
If all inputs are dense, the non-distributed form is more efficient because it results in fewer terms and has the same computational cost per term. However, if $X_{ij}$ is sparse and $U_{i},V_{j}$ are dense, then the distributed form is more efficient because all terms can be computed in time linear w.r.t. the sparsity of $X_{ij}$. Note that the squaring operation here is a pointwise function, not a matrix multiplication.
To take advantage of this potentially asymptotic performance improvement, Galley performs a greedy search for the optimally distributed expression. At each step, it considers all single applications of distributivity in the expression. It then runs variable elimination for each (described later in this section) and computes the cost an optimal logical plan. If applying distributivity improved on the cost of the original expression, it continues. If not, it returns the optimal logical plan discovered so far. Lastly, we additionally consider the expression derived from applying distributivity exhaustively.
5.2. Cost Model
Overall, Galley’s logical optimizer attempts to minimize the time required to execute the logical program. Because logical queries do not correspond to concrete implementations, our logical cost model aims to approximate this time without reference to the particular implementation that the physical optimizer will eventually decide on. This approximation considers two factors: (1) the number of non-fill entries in the output tensor and (2) the amount of computation (i.e., the number of FLOPs) needed to produce the output. The former corresponds to the size of the tensor represented by Agg, $nnz(\texttt{Agg})$, and the latter corresponds to the tensor size represented by the MapExpr within, $nnz(\texttt{MapExpr})$. We assume that the inputs are in memory; hence, there is no cost for reading inputs from disk. We then perform a simple regression to associate each cost with a constant, and we add them to produce our overall cost, $c$, as follows:
$cost\approx a*nnz(\texttt{Agg})+b*nnz(\texttt{MapExpr})$
To estimate $nnz(\texttt{Agg})$ and $nnz(\texttt{MapExpr})$, we use the sparsity estimation framework described in Sec. 7.
5.3. Variable Elimination
The core of our logical optimizer is the variable elimination (VE) framework. With this view, the logical plan for an input query is defined by an order on the indices being aggregated over, i.e., an elimination order. Given this order, we can construct a valid logical plan by iterating through the elimination order one index at a time in order to (1) identify the minimal sub-expression needed to aggregate over it, (2) create a new logical query representing the result of that sub-expression, and (3) replace it in the original query with an alias to the result. At the end of this process, the remaining query no longer requires any aggregation and therefore is itself a logical query.
Example 0.
Consider optimizing the following matrix chain multiplication:
$E_{im}=\sum_{jkl}A_{ij}B_{jk}C_{kl}D_{lm}$
The elimination order $jkl$ corresponds to a left-to-right multiplication strategy because eliminating $j$ from the expression first requires performing the matrix multiplication between $A$ and $B$. Eliminating $k$ then requires multiplying that intermediate result with $C$, and so on. Concretely, this produces the following sequence of logical queries:
    Query(I1, Agg(+, j, Map(*, A[i,j], B[j,k])))
    Query(I2, Agg(+, k, Map(*, I1[i,k], C[k,l])))
    Query(E, Agg(+, l, Map(*, I2[i,l], D[l,m])))
Similarly, the elimination order $lkj$ corresponds to a right-to-left strategy, and the order $klj$ to a middle-first strategy.
Unlike traditional VE for sum-product queries, we support complex trees of pointwise operators and aggregates. This makes identifying minimal sub-queries challenging since we must carefully examine the expression’s algebraic properties. Given a strategy for this, the core problem of optimizing VE is to search the space of elimination orders for the most efficient one. In the worst case, this takes exponential time w.r.t. the number of indices being aggregated over. In the following sections, we describe how we identify minimal sub-queries and our search algorithm for finding the optimal elimination order.
5.4. Identifying Minimal Sub-Expressions
We now explain how to identify the minimal sub-expressions (MSEs) needed to eliminate an index. In sum-product expressions, the MSE is simply a summation over that index wrapping the product of the tensors that are indexed by it. However, it is not immediately clear how to do this for more complex input programs. Fortunately, we show that identifying MSEs corresponds to a careful traversal down the annotated expression tree, examining the algebraic properties of the operation at each node to determine how to proceed.
Figure 5. Annotated expression tree for logistic regression
$\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
Annotated Expression Tree. The annotated expression tree (AET) is constructed by examining the nested structure of Agg, Map, Input, and Alias nodes in the input query. To do this, Galley first removes all Agg nodes and annotates their inner expressions with (Idx, Op). It then replaces all Map nodes with their operator to get the final tree, where every internal node is a pointwise function and every leaf is either an Input or an Alias.
Example 0.
Fig. 5 shows the annotated expression tree for our logistic regression example after distributing the multiplication as in Eq. (2). The sigmoid function is the outermost layer of the expression, so it appears at the top of the tree. The summations all occur just inside the sigmoid function, so they annotate the top multiplication operator, as denoted by the red arrow.
Given the AET, Galley identifies an index’s MSEs by starting at the node where it is annotated and traversing downwards according to the algebraic properties of each internal node. We now describe the traversal rules for functions that are distributive, non-distributive, and commutative with respect to the aggregation operator.
Distributive Functions. When we reach a function that distributes over the aggregate (e.g., * and $\sum$), we examine how many of the children, subtrees of the AET, contain the current index. If one child contains the index, we traverse down that child’s branch, i.e., we factor the other children out of the aggregate. If multiple children contain the index, we cannot go any further down the tree, so we wrap the sub-tree rooted at that node in the aggregate and return it as our MSE. If the function is commutative and associative, we are slightly more precise and include only the children that contain the index.
Commutative, Identical Functions. When the node’s function is the same as the aggregate function and is commutative, we can push the aggregate down to each child independently. For example, we can transform the expression $\sum_{i}A_{i}+B_{i}$ into $\sum_{i}A_{i}+\sum_{i}B_{i}$. For all children that contain the index, we add the result of traversing down its branch to the list of MSEs and replace it with an alias to the result. If a child does not contain the index, then we need to account for the repeated application of the aggregate function. To do this, we require a function $g(x,n)=f(x,...,x)$ that represents the repeated application of our aggregate function $f$. When $f$ is addition, $g$ is multiplication. When $f$ is idempotent, $g(x,n)=x$. We then wrap each non-index child in a Map with function $g$ and the size of the index’s dimension as the second child. Further, we can choose to push down along some of these dimensions and not others, e.g. $\sum_{i}(A_{i}+B_{i}+C_{i})=\sum_{i}(A_{i}+B_{i})+\sum_{i}(C_{i})$. To allow this, we represent each path that a variable can be pushed down with a virtual index, and we include these in our elimination ordering. This increases our optimization space but allows us to capture more aggregation strategies.
Blocking Functions. A function that does not distribute or commute with our aggregate function is called a blocking function. When we reach a blocking function in our traversal, we simply wrap it in our aggregate and return the sub-tree as an MSE. For example, the expression $\sum_{j}\sqrt{A_{ij}B_{jk}}$ cannot be rewritten as $\sqrt{\sum_{j}A_{ij}\sum_{j}B_{jk}}$ because $\sqrt{}$ is a blocking function.
Discussion. Galley builds upon and extends the FAQ framework for optimizing conjunctive queries with aggregation(Khamis et al., 2016). This framework explored the optimization of queries with the following form, where each $\bigoplus^{(i)}$ is either equal to or forms a semi-ring with $\otimes$:
$\bigoplus_{v_{1}}^{\,\,\quad(1)}\cdots\bigoplus^{\,\,\quad(k)}_{v_{k}}F^{1}_{V_{1}}\otimes\cdots\otimes F^{k}_{V_{k}}$
Though this framework captures many important problems, it lacks the flexibility needed to support a general tensor processing system. Consider a slightly modified version of the SDDMM kernel:
$\sum_{j}A_{ik}(B_{ij}+C_{jk})$
This expression is not an FAQ query because it mixes addition and multiplication in the pointwise expression. Similarly, our logistic regression example, $\sigma(\sum_{j}X_{ij}\theta_{j})$, cannot be expressed as an FAQ because the aggregate occurs within $\sigma$. Galley extends this framework by accommodating arbitrary pointwise function composition and arbitrary placement of aggregates within expressions.
5.5. Restricted Elimination Orders
Before we can search for the optimal elimination order, we must define the space of valid elimination orders. Depending on the input’s structure, the order in which indices can be eliminated might be restricted. This could be due to non-commutative aggregates or aggregate placement. The former occurs when an aggregate wraps another aggregate it does not commute with. For example, if the expression is $\max_{i}\sum_{j}A_{ij}$, we must perform the summation before handling the maximum because max and $\sum$ do not commute. The placement issue arises when an aggregate wraps another aggregate but cannot reach it via the traversal described above, e.g., $\sum_{i}\sqrt{\sum_{j}A_{ij}}$; in this case, the inner aggregate must be performed first. Collectively, these restrictions form a partial ordering on the index variables that must be respected when we enumerate elimination orders.
5.6. Search Algorithms
With the VE approach, we have simplified the complicated issue of high-level optimization to the discrete problem of choosing an optimal order on the aggregated index variables. We start by revisiting our example from Fig. 5. The input query is the following,
    Query(X, Map($\sigma$,
                Agg(+,p,c,j,
                    Map(*,S[i,p,c],
                        Map(+,
                            Map(*, P[p,j], $\theta$[j]),
                            Map(*, C[c,j], $\theta$[j]))))))
The elimination order for this expression is an ordering of the indices $\{p,c,j^{1},j^{2}\}$. Note that $j^{1}$ and $j^{2}$ are virtual indices, introduced to represent the two paths that $j$ can be pushed down in the AET due to the addition operator. Galley’s logical optimizer searches through these possible orders to find the most efficient one. In this case, it would choose $[j^{1},j^{2},p,c]$, resulting in the following logical plan,
    Query(A1, Agg(+, j, Map(*, P[p,j], $\theta$[j])))
    Query(A2, Agg(+, j, Map(*, C[c,j], $\theta$[j])))
    Query(A3, Agg(+, p, c, Map(*, S[i,p,c],
                                Map(+, A1[p], A2[c]))))
    Query(X, Map($\sigma$, A3[i]))
We now present two algorithms to search for that optimal order using the tools described above.
Greedy. The greedy approach simply chooses the cheapest index to aggregate at each point by finding the minimal sub-query for each index and computing its cost. The cheapest index’s minimal sub-query is removed from the expression and appended to the logical plan; it is then replaced in the remaining query with an alias to the result. This continues until no aggregates remain in the expression.
Branch-and-Bound. The branch-and-bound approach computes the optimal variable order and occurs in two steps. The first step uses the greedy algorithm to produce an upper bound on the cost of the overall plan; the second performs a dynamic programming algorithm. In the dynamic programming step, the keys of the memo table are unordered sets of indices, and the values are a tuple of partial elimination orders, residual queries, and costs. The algorithm initializes the table with the empty set and a cost of zero. At each step, it iterates through table entries and attempts to aggregate out another index. It then uses the cost bound from the first step to prune entries from the memo table whose cost exceeds the bound; doing so is valid because costs monotonically increase as more indices are added to the set. At the end of this step, the algorithm returns the index order associated with the full set of indices.
6. Physical Optimizer
Each query in the logical dialect roughly corresponds to a single loop nest and materialized intermediate. However, several decisions remain about how the kernel is computed, including: (1) the loop order over the indices, (2) the format (i.e., layout) of the result, and (3) the protocol (i.e., search algorithm) for accessing each index of each input. The physical optimizer makes these decisions.
6.1. Loop Order
The loop order of a tensor kernel determines the order of access to the inputs. An optimal order results in early pruning of iterations due to early intersection of sparse inputs. Intuitively, this is similar to selecting a variable order for a worst-case optimal join algorithm. To find the optimal order, Galley has a cost model for loop orders based on sparsity estimates and an optimization algorithm that uses this cost model to guide its search.
Cost Model
The cost of a loop order equals the number of iterations each level of the loop nest incurs.
Example 0.
Consider matrix chain multiplication over three sparse matrices, $A$, $B$, and $C$, where
(3) $D[il]=\sum_{jk}A[ij]*B[jk]*C[kl]$
Suppose that $A$ has only a single non-zero entry and that $B$ and $C$ have 5 non-zero entries per column and per row. In this case, the loop order $ijkl$ is significantly more efficient than $lkji$. In the former, the first two loops, over $i$ and $j$, incur only a single iteration because they are bounded by the size of $A$. The third and fourth incur $5$ and $5^{2}$ iterations, respectively, because there are only $5$ non-zero $k$’s per $j$ in $B$ and $5$ non-zero $l$’s per $k$ in $C$. In the latter, the first two loops iterate over the full matrix $C$ despite most of those iterations not leading to useful computation.
The last piece of our cost model is the cost of transposition. If an input’s index order is not concordant with the loop order, it must be transposed before the query can be executed. This imposes a cost that is linear in that input’s size. We ignore the cost of transposing intermediates because we they are always materialized in a concordant order.
Formally, let $Q$ be the pointwise expression in our kernel, and let $Q_{(i_{1},\ldots,i_{k})}$ be the restriction of that expression to just the sequence of index variables $i_{1},\ldots,i_{k}$. Further, let $\mathbf{A}^{(i_{1},\ldots,i_{k})}$ be the set of input tensors which are not concordant with $i_{1},\ldots,i_{k}$. Then, we can define the cost of a loop order as follows,
$cost(Q,(i_{1},\ldots,i_{k}))\approx\sum_{j=1}^{k}nnz(Q^{(i_{1},\ldots,i_{j})})+\sum_{A\in\mathbf{A}^{(i_{1},\ldots,i_{k})}}|A|$
In practice, we further refine this model to take into account the number and kind of tensor accesses at each level of the loop nest.
Optimization Algorithm.
To optimize the loop order, we combine this cost model with a branch-and-bound, dynamic programming algorithm. In the first pass, the optimization algorithm selects the cheapest loop index at each step until reaching a full loop order. This produces an upper bound on the optimal execution cost, which the algorithm uses to prune loop orders in the second step. This step applies a dynamic programming algorithm. Taking inspiration from Selinger’s algorithm for join ordering, each key in the DP table is a set of index variables and a set of inputs. The former represent the loops that have been iterated so far, and the latter represents a set of inputs that must be transposed.
6.2. Intermediate Formats
Once the loop order has been determined, the physical optimizer selects the optimal format for each query’s output. First, Galley selects the order of the indices in the output to be concordant with either the loop order of the kernel where it will be consumed or the order requested by the user. It then selects a format for each index of the output (e.g., dense vector, hash table, etc.). Two factors affect this decision: (1) the kind of writes being performed (sequential vs random) and (2) the sparsity of the tensor at this index. The former is important because many formats (e.g., sorted list formats) permit only sequential construction. These formats can only be used if the output indices up to that point form a prefix of the loop order.
When considering sparsity, Galley balances the fact that denser formats tend to be more efficient due to their memory locality and simplicity, while sparser formats are asymptotically smaller for sparse outputs. To describe this trade-off, we hand selected sparsity cutoffs between fully sparse, bytemap, and fully dense formats. To determine a particular output index’s format, the physical optimizer first determines the sparsity at this index level and uses our cutoffs to determine which category of formats to consider. Then, it checks whether sequential or random writes are being performed and selects the most efficient format that supports the write pattern.
6.3. Merge Algorithms
The final decision the physical optimizer makes concerns the algorithm it will use to perform each loop’s intersection. While there are more complex strategies, we adopt instead a minimal approach and select a single input to iterate over for each loop. The physical optimizer then probes into the remaining inputs. It makes this selection by estimating the number of non-zero indices that each input has, conditioned on the indices in the outer loops. This resembles the approach taken in (Wang et al., 2024) for optimizing WCOJ.
6.4. Common Sub-Expression Elimination
Galley takes a straightforward approach to avoiding redundant computation. Once a physical plan has been generated, the right hand side of each physical query is canonicalized and hashed. When two physical queries result in the same hash, the latter query is removed from the plan and all references to it are replaced with a reference to the result of the former. This is helpful for caching small computations like transpositions, but it is also useful for reducing the overhead of applying distributivity which often results in duplicate sub-expressions.
7. Sparsity Estimation
We now describe how Galley performs the sparsity estimation that guides our logical and physical optimizers. First, we explore the subtle correspondence between sparsity and cardinality estimation. We then present a minimal interface for sparsity estimation inspired by this correspondence, after which we examine two implementations of this framework, i.e., the uniform estimator and the chain bound.
7.1. Sparsity and Cardinality Estimation
Sparsity estimation is highly related to cardinality estimation in databases. However, translating methods for the latter to the former requires analyzing the algebraic properties of our tensor programs. For example, let $A_{ij}$ and $B_{jk}$ be sparse matrices with a fill value of 00, and let $R_{A}(I,J)$ and $R_{B}(J,K)$ be relations that store the indices of their non-zero entries. Assume we are performing the following,
$C_{ijk}=A_{ij}B_{jk}$
In this case, the number of non-zero values in $C$ is precisely equal to the size of the conjunctive query
$nnz(C)=|R_{A}(I,J)\bowtie R_{B}(J,K)|$
The correspondence results from the fact that 00 is the annihilator of multiplication (i.e., $x*0=0\forall x$), so any non-zero entry $ijk$ in the output must correspond to a non-zero $ij$ in $A$ and a non-zero $jk$ in $B$. Consider the following instead:
$C_{ijk}=A_{ij}+B_{jk}$
In this case, a nonzero $ijk$ in the output can result from a non-zero $ij$ in $A$ or a non-zero $jk$ in $B$. In traditional relational algebra, where relations are over infinite domains, this kind of disjunction would result in an infinite relation. However, tensors have finite dimensions, so we can introduce relations that represent the finite domains of each index, e.g., $D_{i}=\{1,...,n_{i}\}$. This lets us represent the index relation of the output as
$nnz(C)=|(R_{A}(I,J)\bowtie D_{k}(K))\cup(D_{i}(I)\bowtie R_{B}(J,K))|$
Finally, we can translate aggregations to the tensor setting as projection operations. Given the statement
$C_{ik}=\sum_{j}A_{ijk}$
we can express the non-zeros entries of $C$ as
$nnz(C)=|\pi_{I,K}(R_{A}(I,J,K))|$
7.2. The Sparsity Statistics Interface
We use our statistics interface to annotate an expression with stats objects at every node of the AST in a bottom-up fashion. Each stats object then represents the sparsity patterns of the intermediate tensor output from that node. Surprisingly, to support sparsity estimation over the varied workloads and arbitrary operators of tensor algebra, we need to implement only a few core functions: (1) a constructor, which produces statistics from a materialized tensor for Input and Alias nodes, (2) a function for annihilating Map nodes (i.e., those whose children’s fill values are the annihilator of its pointwise function), which merges the children’s statistics, (3) a function for non-annihilating Map nodes, which merges the children’s statistics, (4) a function for Agg, which adjusts the input’s statistics to reflect an aggregation over some set of indices, and (5) an estimation procedure, which estimates the number of non-fill entries based on statistics about a tensor.
7.3. Supported Sparsity Estimators
7.3.1. Uniform Estimator
The simplest statistic that can be kept about a tensor is the number of non-fill (e.g., non-zero) entries. The uniform estimator uses only this statistic and relies on the assumption that these entries are uniformly distributed across the dimension space. This corresponds to System-R’s cardinality estimator with the added assumption that the size of the index attribute’s active domain equals the size of the dimension (Selinger et al., 1979).
Constructor
Given a tensor $A_{i_{1},\ldots,i_{k}}\in\mathbf{R}^{n_{i_{1}}\times\ldots\times n_{i_{k}}}$, this function simply counts the non-fill values in the tensor, $nnz(A)$, and notes the dimension sizes $n_{i_{1}},\ldots,n_{i_{k}}$.
Map (Annihilating)
To handle an annihilating pointwise operation, this function calculates the probability that a point in the output was non-fill in all inputs, then multiplies this by the dimension space of the output. For a set of inputs $A^{(1)}_{I_{1}}\ldots A^{(l)}_{I_{l}}$ and output $C_{I_{C}}$, where each $I_{j}$ is a set of indices, this probability is
$nnz(C)\approx\left(\prod_{i\in I_{C}}n_{i}\right)\cdot\left(\prod_{j}\frac{nnz(A_{j})}{\prod_{i\in I_{j}}n_{i}}\right)$
Map (Non-Annihilating)
To handle an non-annihilating pointwise operation, this function calculates the probability that an entry in the output was fill in all inputs. Then, it takes the compliment to get the probability that it was non-fill in all inputs and multiplies this by the output dimension space. Using the preceding notation:
$nnz(C)\approx\left(\prod_{i\in I_{C}}n_{i}\right)\cdot\left(1-\prod_{j}\left(1-\frac{nnz(A_{j})}{\prod_{i\in I_{j}}n_{i}}\right)\right)$
Aggregate
Given an input tensor $A_{I}$ to aggregate over the indices $I^{\prime}$, this function computes the probability that an output entry is non-fill by calculating the probability that at least one entry in the subspace of the input tensor was not fill:
$nnz(C)\approx\left(\prod_{i\in I\setminus I^{\prime}}n_{i}\right)\cdot\left(1-\left(1-\frac{nnz(A_{I})}{\prod_{i\in I}n_{i}}\right)^{\prod_{i\in I^{\prime}}n_{i}}\right).$
Estimate
The estimation function simply returns the current tensor’s stored cardinality statistic.
7.3.2. Degree Statistics and the Chain Bound
Galley stores degree statistics as the default, and it uses them to compute upper bounds on the number of non-fill entries in intermediate expressions. A degree statistic, denoted as $D_{A}(X|Y)$, stores the maximum number of non-fill entries in the $X$ dimensions conditioned on the $Y$ dimensions for a tensor $A$. For example, given a matrix $A_{ij}$, then $D_{A}(i|j)$ is the maximum number of non-fill entries per column, and $D_{A}(ij|\emptyset)$ is the total number of non-fill entries in the matrix. This approach follows work in cardinality bounding that has been shown to produce efficient query plans in the relational setting (Deeds et al., 2023; Hertzschuch et al., 2021; Cai et al., 2019).
Constructor
This function first computes the boolean tensor representing the input’s sparsity pattern. Then, to calculate each degree statistic, it sums over the $X$ dimensions and takes the maximum over the $Y$ dimensions. The set of degree statistics for a tensor $A_{I}$ is denoted $\mathcal{D}_{A_{I}}$.
Map (Annihilating)
Annihilating map operations serve as conjunctive queries with respect to the sparsity patterns of the inputs. Therefore, any degree statistics that are valid for an input are also valid about the output. This function computes the statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by a union
$\mathcal{D}_{C}=\bigcup_{j}A^{(j)}_{I_{j}}$
Map (Non-Annihilating)
In this case, the function must be especially careful to ensure that it maintains the upper bounds. First, it extends the degree constraints from each input to cover the full set of indices. For example, if we have $DC_{X}(i|j)$ and want to extend it to the dimension $k$, the function computes $D_{X}(ik|j)=DC(i|j)*n_{k}$. Then, it computes degree statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by addition:
$D_{C}(X|Y)=\sum_{j}D_{A^{(j)}}(X|Y)$
Estimator
This function calculates an upper bound (eq. performs sparsity estimation) using the breadth-first search approach described in (Chen et al., 2023). Intuitively, each set of indices forms a node in the graph, and each degree constraint is a weighted edge from $Y$ to $X$. Its search begins with the empty set; it then uses a breadth-first search to find the shortest weighted path to the full set of indices $I$. The product of the weights along this path bounds the number of non-zeros in the result.
Figure 6. ML Inference Over Joins
8. Experimental Evaluation
In this section, we evaluate the effectiveness of our optimizer on a variety of workloads: (1) ML algorithms over joins, (2) subgraph counting, and (3) breadth-first search. We choose those workloads because they exercise different aspects of our optimizer on real-world use-cases: ML algorithms over joins require careful logical optimizations over programs with mixtures of dense and sparse inputs and non-linear operators; subgraph counting requires both logical and physical optimization of complex sum-product expressions over highly sparse inputs and demonstrates Galley’s advantage over a relational engine even for very sparse workloads; breadth-first search requires careful selection of tensor formats over the course of the computation, showing the benefit of physical optimization for even simple computations. Compared to hand-optimized solutions and alternative approaches, Galley is highly computationally efficient while requiring only a concise, declarative input program from the user. Overall, we show that Galley:
[…]
  • Selects optimal formats for intermediates, outperforming both fully dense and sparse formats for 4/5 graphs in a BFS application.
Experiment Setup
These experiments are run on a server with an AMD EPYC 7443P Processor and 256 GB of memory. We implemented Galley in the programming language Julia, and the code is available at https://anonymous.4open.science/r/Galley-21BF/. We used the sparse tensor compiler Finch [^2] https://github.com/FinchTensor/Finch.jl for execution, and all experiments are executed using a single thread. Unless otherwise stated, Galley uses the chain bound described in Sec. 7.3.2 for sparsity estimation. Experiments for all methods are run five times, and the mean execution time is reported. We perform all experiments on a warm cache, and we separately report the compilation and optimization times.
8.1. Machine Learning Algorithms
To explore end-to-end program optimization, we experiment with simple ML algorithms over joins, represented entirely in tensor algebra. For this, we use the TPC-H benchmark at a scale factor of .25.25.25.25 in order to keep the comparison methods in memory [^3] Galley’s execution tends to be less memory-intensive, and we have successfully run it at scale factor 1.5.. We consider two join queries on this dataset: star and self-join. First, we perform a star join over the line items table to gather features about suppliers, parts, orders, and customers. This is expressed as follows, where $L,S,P,O,$ and $C$ are tensors representing the line items, suppliers, parts, orders, and customers tables, respectively:
$X_{ij}=\sum_{spoc}L_{ispoc}(S_{sj}+P_{pj}+O_{oj}+C_{cj})$
The non-zero values in $S,P,O$ and $C$ are disjoint along the $j$ axis, so the addition in this expression serves to concatenate features from each source, resulting in 139 features after one-hot encoding categorical features. The self-join query compares line items for the same part based on part and supplier features. In this case, the feature data is a 3D tensor because the data points are keyed by pairs of line items:
$X_{i_{1}i_{2}j}=\sum_{s_{1}s_{2}p}L_{i_{1}s_{1}p}L_{i_{2}s_{2}p}(S_{s_{1}j}+S_{s_{2}j}+P_{pj})$
Given these definitions, we consider a range of ML algorithms: (1) linear regression inference, (2) logistic regression inference, (3) covariance matrix calculation, and (4) neural network inference. For comparison, we also implement two versions of each of these using the Finch compiler. The dense version uses a fully dense format to represent the features, whereas the sparse version uses a sparse level for the features to compress the one-hot encoding. We also implemented these algorithms using the Pandas and Numpy libraries for reference.
These algorithms stress the ability of Galley to handle both combinations of complex operators and combinations of sparse and dense inputs. The definitions of the feature tensors combine pointwise multiplication and addition, and algorithms like logistic regression and neural networks wrap these definitions in more non-linear operators (e.g. relu and sigmoid) and aggregates. Further, while the line item tensor is highly sparse, both the feature and parameter tensors are moderately to fully dense. In all of these examples, Galley is able to manipulate these complex expressions to avoid materializing large, sparse intermediates. Instead, its logical optimizer aggressively pushes the computation and aggregation down to the smaller, denser inputs. Consider logistic regression over the star query feature matrix. In this case, Galley distributes the parameter vector over the addition and performs the summations $\sum_{j}S_{sj}\theta_{j},\sum_{j}P_{pj}\theta_{j},\sum_{j}O_{oj}\theta_{j},$ and $\sum_{j}C_{cj}\theta_{j}$ before handling the remaining operation in one kernel. By doing this, it materializes a dense vector at every point and removes the feature dimension before operating on the large, sparse line item tensor.
Execution Time. Fig. 6 shows that the execution time of Galley’s optimized programs is .5300×.5-300$\times$.5 - 300 × faster than the sparse Finch implementation. For the regression/neural network problems, this stems from pushing the parameter vector/matrix down to the feature matrices as discussed above. For the covariance calculation over the self join, Galley fully distributes the multiplication over the addition. Each distributed term is the product of four copies of the lineitem table $l_{i_{1},s_{1},p_{1}},l_{i_{2},s_{2},p_{1}}l_{i_{1},s_{3},p_{2}}l_{i_{2},s_{4},p_{2}}$ and one feature matrix from each $X$, e.g. $S_{s_{1},j}P_{p_{1},k}$. Within each term, Galley performs the summation over $i_{1}$ and $i_{2}$ first, before handling the summation of $p_{1}$ or $p_{2}$. Because $i_{1}$ and $i_{2}$ are keys, unlike $p_{1}$ and $p_{2}$, these initial sums produce small, dense intermediates, speeding up the whole computation. Lastly, the range of Galley’s execution times for the five iterations of each experiment was never more than $25\%$ of the mean.
Optimization Time. Fig.  6 also shows that Galley’s optimizer has a reasonable overhead in this setting. For all experiments except for Covariance (SQ), Galley is faster than competing methods even when including the optimization time. Concretely, optimization takes .$15-3.0$ seconds on these workloads. The covariance calculations take longer to optimize due to the number of terms produced when they are fully distributed.
Figure 7. Subgraph Counting Execution Time
Figure 8. Subgraph Counting Optimization Time
Figure 9. Subgraph Counting Compilation Time
Figure 10. Sparsity Estimator Comparison
8.2. Subgraph Counting
In this section, we test Galley’s ability to optimize programs with a large number of highly sparse inputs. To do this, we implement a few common sub-graph counting benchmarks. The conversion from sub-graph counting to sparse tensor algebra is straightforward and results in traditional einsum expressions. Suppose you are counting the occurrences of $H(V,E)$ in a data graph $G$ with adjacency matrix $M$; we can represent the count as
$c=\sum_{v_{i}\in V}\prod_{(v_{i},v_{j})\in E}M_{v_{i}v_{j}}$
To handle vertex labels, we add sparse binary vector factors for each labeled vertex. We use subgraph workloads from the G-Care benchmark and the paper ”In-Memory Subgraph: an In-Depth Study”(Sun and Luo, 2020; Park et al., 2020). We restrict the latter benchmark to query graphs with up to 8 vertices (hence the ”_lite” suffix). Because this is a relational workload, we compare it with DuckDB, a modern OLAP database (Raasveldt and Mühleisen, 2019). To separately discern the impact of logical vs physical optimization and our use of Finch, we provide a version of Galley that executes each logical query with a SQL query run on DuckDB. Finally, we provide results for the greedy logical optimizer, as well.
Logical Optimization. Fig. 7 shows the execution time for each method and benchmark. The comparison between ‘DuckDB‘ and ‘Galley + DuckDB Backend‘ demonstrates the benefits of Galley’s logical optimizer. Galley’s logical optimizer breaks down the program into a series of aggregations which effectively minimize the necessary computation and materialization, resulting in a dramatic performance improvement. This has the largest impact on graphs with high degree values and/or high skew like the social network graphs, ‘dblp_lite‘ and ‘youtube_lite‘. In these cases, not pushing down aggregates can result in very large intermediate results. On low-degree graphs like ‘aids‘, these optimizations have less impact. Further, DuckDB hits the 300 second timeout on 50 out of 120 queries in the youtube_lite benchmark. In contrast, Galley never times out across all workloads.
Physical Optimization. The impact of Galley’s physical optimizer can be seen by comparing ‘Galley‘ with ‘Galley + DuckDB Backend‘. By producing efficient STC kernels, Galley’s median execution is up to 8x faster than DuckDB even with the same logical plan. This shows that Galley is selecting efficient loop orders and formats which allows it to take advantage of the highly flexible language of STCs.
Optimization Time. Fig. 8 shows the mean optimization time for each method on each workload. Galley has a mean optimization time of less than .15.15.15.15 seconds across all workloads, approaching the time taken by the highly efficient DuckDB optimizer.
Compilation Time. Because it performs compilation using Finch at runtime, Galley incurs a compilation latency the first time that it invokes each unique STC kernel. Fortunately, these kernels are automatically cached by Finch, reducing this cost when workloads repeatedly use similar kernels. We show the mean compilation time for each subgraph workload in Fig. 9. On the simpler workloads, which often reuse kernels, this cost is minimal. However, the more complex workloads both reuse kernels less and require compiling more complex kernels, significantly increasing compilation time. This suggests a need for faster STC compilation or the use of an interpreted engine for complex queries.
Sparsity Estimation. Finally, in Fig. 10, we use the sub-graph matching workloads to compare sparsity estimators and their effect on performance. Across all workloads, we see that the chain bound significantly out-performs the uniform estimator. This is because graph datasets tend to be skewed and correlated. Vertices have drastically different edge counts, and vertices with many edges tend to connect to vertices with many edges. This violates the implicit assumptions of the uniform estimator, leading it to optimistically materialize large intermediates. Due to this, the uniform estimator runs out of memory on many queries in the ‘youtube_lite‘ workload, so it is not pictured in this figure.
Figure 11. BFS Execution Time
8.3. Breadth-First Search
To demonstrate the importance of selecting optimal physical formats for tensors, we implement a simple breadth-first search algorithm using Galley and different hand-coded Finch implementations. Both systems receive a single iteration at a time, and the total execution time across all iterations is reported. The core computation for each iteration is a masked sparse matrix times sparse vector operation that computes the new frontier vector. The main optimization decision is the visited and frontier vectors’ formats. The former grows monotonically over the course of the algorithm, while the number of non-zeros in the latter forms a curve, peaking in the middle iterations. We provide two implementations of Finch, using either a sparse or a dense vector for both intermediates. Fig. 11 shows that Galley is significantly faster than both dense and sparse approaches for 4 of the 5 graphs and is competitive for all graphs. During a single run of BFS, Galley begins with a sparse visited vector before switching to bytemap and dense formats as more nodes are discovered. The frontier vector begins as sparse, becomes dense, then again becomes sparse for the last few iterations as most nodes are deactivated. For 4/5 graphs, the total optimization time for Galley (not depicted in the figure) is less than .25.25.25.25 seconds. RoadNet has a very wide diameter leading to many iterations and a total optimization time of $4$ seconds. This experiment demonstrates the utility of sparsity-aware format selection, and future work should consider ways to amortize optimization time for iterative workloads.
9. Conclusion and Limitations
This paper introduced Galley, a system for efficient declarative sparse tensor programming. We describe and then demonstrate how it optimizes high-level program structure with its logical optimizer and how it lowers that program to an efficient implementation with its physical optimizer. Guiding these decisions are sparsity estimates of intermediate expressions, and we show that these estimates can be computed for arbitrary tensor algebra programs by implementing a minimal 5-function interface. Finally, we present results from our evaluation of experiments with Galley on ML algorithms over structured feature data, sub-graph counting, and breadth-first search.
We are excited to enrich Galley further with new optimizations in the future. Currently, Galley lacks support for complex loop structures (e.g., a single outer FOR loop that wraps multiple inner FOR loops) and parallelism. Both areas could benefit from cost-based optimization. Similarly, Galley does not consider hard memory constraints during optimization, but our use of cardinality bound methods provides an avenue for addressing this in future work.
Acknowledgements.
This work was supported by NSF SHF 2312195, NSF IIS 2314527, and DARPA SBIR: HR001123C0139.